# Parallel transport in rotating frames and projective holonomic quantum computation

### v3

@@ II   Conventional nonadiabatic holonomic quantum computation
Holonomic quantum computing uses holonomy to implement gates. Holonomy refers to the geometric phenomenon that when the computational space is transported along a closed path, a co-transported computational frame, that is, a frame that spans the computational space, may end up rotated relative to its original configuration, even though the transport is parallel, that is, performed without causing any time-local rotation of the computational frame.
A holonomic gate is a unitary operator representing the rotation of a computational frame resulting from a cyclic parallel transport of the computational space. A more precise description of a holonomic gate is as follows (see also [1, 2, 3, 21]). Consider a quantum system with a Hilbert space $\mathcal{H}$ and a computational subspace $\mathcal{R}$ [^1] Thus, computational input states are prepared to have support in $\mathcal{R}$ and are manipulated by unitary operators on $\mathcal{R}$, a.k.a gates. Let $\Pi_{t}$ be a one-parameter family of isometric embeddings of $\mathcal{R}$ in $\mathcal{H}$ that transport $\mathcal{R}$ in a loop:
@@ II.B   The isoholonomic inequality and an execution time estimate for holonomic gates
$I(H_{t};\mathcal{R}_{t})=-\frac{1}{2}\operatorname{tr}\big{(}[H_{t},P_{t}]^{2}\big{)}$ (9)
@@ II.C   Parallel transport in a rotating frame
$|\psi_{t}^{\textsc{rf}}\rangle$ $=R_{t}|\psi_{t}\rangle,$ (11)
$O^{\textsc{rf}}_{t}$ $=R_{t}O_{t}R^{\dagger}_{t},$ (12)
$H^{\textsc{rf}}_{t}$ $=R_{t}H_{t}R^{\dagger}_{t}+i\dot{R}_{t}R_{t}^{\dagger}.$ (13)
@@
The computational space being transported in a loop can also be formulated more invariantly: Let $P$ be the orthogonal projection onto $\mathcal{R}$. That $\mathcal{R}_{t}$ is closed can be expressed as $\langle\psi_{\tau}|P|\psi_{\tau}\rangle=1$ for all unit vectors $|\psi\rangle$ in $\mathcal{R}$. In the rotating frame picture, observables evolve over time according to Eq. (12), ensuring that all statistical quantities remain invariant under the change of reference frame. Consequently, $\mathcal{R}_{t}$ is closed if and only if $\langle\psi_{\tau}^{\textsc{rf}}|P_{\tau}^{\textsc{rf}}|\psi_{\tau}^{\textsc{rf}}\rangle=1$ for all unit vectors $|\psi^{\textsc{rf}}\rangle$ in $\mathcal{R}^{\textsc{rf}}$.
@@ III.C   Projective holonomic gates and projectively parallel transporting Hamiltonians
A Hamiltonian $H_{t}$ with time translation operator $U_{t}$ is said to be projectively parallel translating if, for some (and hence any) $n$-frame $V$ for $\mathcal{R}$, the curve $V_{t}=U_{t}(V)$ satisfies $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$ and thus represents a projective horizontal curve. Writing $|v_{k}\rangle$ for the $k$th component of $V$ and $|v_{k;t}\rangle$ for $U_{t}|v_{k}\rangle$, the projective parallel transport condition becomes
@@
for some real-valued $\epsilon_{t}$; c.f. [9, Eq. (25)]. Unlike the conventional parallel transport condition (3), the projective parallel transport condition is gauge invariant in the sense that if $H_{t}$ is projectively parallel transporting, so is $H_{t}+\epsilon_{t}\mathbbm{1}$ for any real-valued function $\epsilon_{t}$.
@@ IV   Tight implementations
The summands $H_{k;t}$ are effective qubit Hamiltonians acting on pairwise perpendicular $2$-dimensional subspaces $\mathcal{E}_{k}$, each spanned by a normalized eigenvector $|v_{k}\rangle$ of $\Gamma$,
@@
and a vector perpendicular to the computational space. [^10] Thus $\mathcal{H}$ is orthogonally decomposed as $\mathcal{H}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}\oplus\cdots\oplus\mathcal{E}_{n}\oplus\mathcal{F}$. The subspace $\mathcal{E}_{k}$ is spanned by an eigenvector of $\Gamma$ and a vector in the orthogonal complement of $\mathcal{R}$. The summand $H_{k;t}$ leaves $\mathcal{F}$ and all the $\mathcal{E}_{l}$’s stationary except $\mathcal{E}_{k}$. As before, we assume that $0\leq\theta_{k}<2\pi$. A key result shown in [22] is that $H_{k;t}$ can be chosen such that
@@
where $B_{k}$ is a Hermitian operator on $\mathcal{E}_{k}$. Choose the eigenbasis of $B_{k}$ such that if $\rho_{k}$ is represented by a Bloch vector $\bm{r}$ and $B_{k}$ by a Rabi vector $\bm{b}$ in $3$-space, then $\bm{r}$ makes an angle $\alpha=\arccos(\theta_{k}/\pi-1)$ with $\bm{b}$; see Figure 2. Adjust the eigenvalues of $B_{k}$ so that $\rho_{k;t}$ becomes periodic with period $\tau$. The curve $\rho_{k;t}$ then has the Fubini-Study speed $(2\pi\theta_{k}-\theta_{k}^{2})^{1/2}/\tau$, and each loop of $\rho_{k;t}$ acquires the geometric phase $\theta_{k}$. However, unless $\theta_{k}=0$ or $\theta_{k}=\pi$, the curve $e^{-itB_{k}}|v_{k}\rangle$ is not horizontal. Therefore, we cannot take $H_{k;t}=B_{k}$.
$\bm{b}$$\bm{r}$$\bm{r}_{t}$$\alpha$
@@
To complete the proof of the existence of $H_{k;t}$, we consider the rotating frame picture obtained by applying the transformation $R_{t}=e^{itB_{k}}$ to the system. In this frame, the parallel transport condition, as given by Eq. (16), takes the form:
@@
We select $H_{k;t}^{\textsc{rf}}$ to be time-independent, $H_{k;t}^{\textsc{rf}}=H_{k}^{\textsc{rf}}$, and choose it such that $|v_{k}^{\textsc{rf}}\rangle=|v_{k}\rangle$ is an eigenvector with eigenvalue $-\langle v_{k}|B_{k}|v_{k}\rangle$. This ensures that the parallel transport condition (40) is satisfied. Also, the evolution equation (39) holds, as $\rho_{k}^{\textsc{rf}}=\rho_{k}$ remains stationary in the rotating frame picture. For more details and a slightly different presentation, see Ref. [22].
$\bm{b}$$\bm{\Omega}^{\textsc{rf}}$$\bm{r}^{\textsc{rf}}$$\bm{b}$$\bm{r}$$\bm{r}_{t}$$\bm{\Omega}$$\bm{\Omega}_{t}$
@@ V   Summary
An oftentimes overlooked fact in nonadiabatic holonomic quantum computation is that the parallel transport condition must be covariantly modified to preserve its physical significance when transitioning between different reference frames. In Section II, we demonstrated how to properly modify the parallel transport condition when shifting from the laboratory frame to a rotating reference frame.
In Section III, we extended the gauge theory of nonadiabatic holonomic quantum computation to a projective gauge theory, enabling holonomic quantum computation with projective gates. This new framework eliminates the problem of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. In addition, we extended the isoholonomic inequality to projective gates and used this extension to derive a sharp estimate for the execution time of projective holonomic gates. This estimate provides a valuable benchmark for developing faster protocols for holonomic quantum gates.
Finally, Section IV reviewed the proof in [22] that if the codimension of the computational space is at least as large as its dimension, then any gate can be tightly implemented using a parallel transporting Hamiltonian. By a tight implementation we mean an implementation that saturates the isoholonomic inequality and, hence, the associated execution time estimate. The proof is constructive and provides guidance in designing schemes for tight implementations of holonomic gates.

### v4

@@ II   Conventional nonadiabatic holonomic quantum computation
Holonomic quantum computing uses holonomy to implement gates. Holonomy refers to the geometric phenomenon that when the computational space is transported along a closed path, a cotransported computational frame, that is, a frame that spans the computational space, may end up rotated relative to its original configuration, even though the transport is parallel, that is, performed without causing any time-local rotation of the computational frame.
A holonomic gate is a unitary operator representing the rotation of a computational frame resulting from a cyclic parallel transport of the computational space. A more precise description of a holonomic gate is as follows (see also [1, 2, 3, 21]). Consider a quantum system with a Hilbert space $\mathcal{H}$ and a computational subspace $\mathcal{R}$ [^1] Thus, computational input states are prepared to have support in $\mathcal{R}$ and are manipulated by unitary operators on $\mathcal{R}$, a.k.a gates. Let $\Pi_{t}$ be a one-parameter family of isometric embeddings of $\mathcal{R}$ in $\mathcal{H}$ that transport $\mathcal{R}$ in a loop in the time $\tau$:
@@ II.B   The isoholonomic inequality and an execution time estimate for holonomic gates
$I(H_{t};\mathcal{R}_{t})=-\tfrac{1}{2}\operatorname{tr}\big{(}[H_{t},P_{t}]^{2}\big{)}$ (9)
@@ II.C   Parallel transport in a rotating frame
$|\psi_{t}^{\textsc{rf}}\rangle$ $=R_{t}|\psi_{t}\rangle,$ (11a)
$O^{\textsc{rf}}_{t}$ $=R_{t}O_{t}R^{\dagger}_{t},$ (11b)
$H^{\textsc{rf}}_{t}$ $=R_{t}H_{t}R^{\dagger}_{t}+i\dot{R}_{t}R_{t}^{\dagger}.$ (11c)
@@
The computational space being transported in a loop can also be formulated more invariantly: Let $P$ be the orthogonal projection onto $\mathcal{R}$. That $\mathcal{R}_{t}$ is closed can be expressed as $\langle\psi_{\tau}|P|\psi_{\tau}\rangle=1$ for all unit vectors $|\psi\rangle$ in $\mathcal{R}$. In the rotating frame picture, observables evolve over time according to Eq. (11b), ensuring that all statistical quantities remain invariant under the change of reference frame. Consequently, $\mathcal{R}_{t}$ is closed if and only if $\langle\psi_{\tau}^{\textsc{rf}}|P_{\tau}^{\textsc{rf}}|\psi_{\tau}^{\textsc{rf}}\rangle=1$ for all unit vectors $|\psi^{\textsc{rf}}\rangle$ in $\mathcal{R}^{\textsc{rf}}$.
@@ III.C   Projective holonomic gates and projectively parallel transporting Hamiltonians
A Hamiltonian $H_{t}$ with time translation operator $U_{t}$ is said to be projectively parallel translating if, for some (and hence any) $n$-frame $V$ for $\mathcal{R}$, the curve $V_{t}=U_{t}V$ satisfies $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$ and thus represents a projective horizontal curve. Writing $|v_{k}\rangle$ for the $k$th component of $V$ and $|v_{k;t}\rangle$ for $U_{t}|v_{k}\rangle$, the projective parallel transport condition becomes
@@
for some real-valued $\epsilon_{t}$; cf. [9, Eq. (25)]. Unlike the conventional parallel transport condition (3), the projective parallel transport condition is gauge invariant in the sense that if $H_{t}$ is projectively parallel transporting, so is $H_{t}+\epsilon_{t}\mathbbm{1}$ for any real-valued function $\epsilon_{t}$.
@@ IV   Tight implementations
The summands $H_{k;t}$ are effective qubit Hamiltonians acting on pairwise perpendicular two-dimensional subspaces $\mathcal{E}_{k}$, each spanned by a normalized eigenvector $|v_{k}\rangle$ of $\Gamma$,
@@
and a vector perpendicular to the computational space. [^10] Thus, $\mathcal{H}$ is orthogonally decomposed as $\mathcal{H}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}\oplus\cdots\oplus\mathcal{E}_{n}\oplus\mathcal{F}$. The subspace $\mathcal{E}_{k}$ is spanned by an eigenvector of $\Gamma$ and a vector in the orthogonal complement of $\mathcal{R}$. The summand $H_{k;t}$ leaves $\mathcal{F}$ and all the $\mathcal{E}_{l}$s stationary except $\mathcal{E}_{k}$. As before, we assume that $0\leq\theta_{k}<2\pi$. A key result shown in [22] is that $H_{k;t}$ can be chosen such that
@@
where $B_{k}$ is a Hermitian operator on $\mathcal{E}_{k}$. Choose the eigenbasis of $B_{k}$ such that if $\rho_{k}$ is represented by a Bloch vector $\bm{r}$ and $B_{k}$ by a Rabi vector $\bm{b}$ in three-space, then $\bm{r}$ makes an angle $\alpha=\arccos(\theta_{k}/\pi-1)$ with $\bm{b}$; see Fig. 2. Adjust the eigenvalues of $B_{k}$ so that $\rho_{k;t}$ becomes periodic with period $\tau$. The curve $\rho_{k;t}$ then has the Fubini-Study speed $(2\pi\theta_{k}-\theta_{k}^{2})^{1/2}/\tau$, and each loop of $\rho_{k;t}$ acquires the geometric phase $\theta_{k}$. However, unless $\theta_{k}=0$ or $\theta_{k}=\pi$, the curve $e^{-itB_{k}}|v_{k}\rangle$ is not horizontal. Therefore, we cannot take $H_{k;t}=B_{k}$.
@@
To complete the proof of the existence of $H_{k;t}$, we consider the rotating frame picture obtained by applying the transformation $R_{t}=e^{itB_{k}}$ to the system. In this frame, the parallel transport condition, as given by Eq. (14), takes the form
@@
We select $H_{k;t}^{\textsc{rf}}$ to be time independent, $H_{k;t}^{\textsc{rf}}=H_{k}^{\textsc{rf}}$, and choose it such that $|v_{k}^{\textsc{rf}}\rangle=|v_{k}\rangle$ is an eigenvector with eigenvalue $-\langle v_{k}|B_{k}|v_{k}\rangle$. This ensures that the parallel transport condition (38) is satisfied. Also, the evolution equation (37) holds, as $\rho_{k}^{\textsc{rf}}=\rho_{k}$ remains stationary in the rotating frame picture. For more details and a slightly different presentation, see Ref. [22].
@@ V   Summary
An often overlooked fact in nonadiabatic holonomic quantum computation is that the parallel transport condition must be covariantly modified to preserve its physical meaning when transitioning between different reference frames. We have shown how to properly modify the parallel transport condition when shifting from the laboratory frame to a rotating reference frame. By observing that central rotations are gauge symmetries, we then naturally extended the gauge theory of nonadiabatic holonomic quantum computation to a projective gauge theory, enabling holonomic quantum computation with projective gates. This new framework eliminates the problem of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. Furthermore, we extended the isoholonomic inequality to projective gates and used this extension to derive a sharp estimate for the execution time of projective holonomic gates, providing a valuable benchmark for developing faster protocols for holonomic quantum gates. We also revisited the proof that if the codimension of the computational space is at least as large as its dimension, then any gate can be tightly implemented using a parallel transporting Hamiltonian. By a tight implementation, we mean an implementation that saturates the isoholonomic inequality and, hence, the associated execution time estimate. The proof is constructive and provides guidance for the design of schemes for tight implementations of holonomic gates.